OpenStudy (toxicsugar22):

Suppose x and y are two possitive numbers. Write an expression for their average, and then for the reciprocal of their average

OpenStudy (anonymous):

what is the average of n numbers?

OpenStudy (toxicsugar22):

i dk am i dupposed to know or is it in the question

OpenStudy (anonymous):

you are supposed to know.

OpenStudy (anonymous):

suppose I had profits of 1000 dollars in one month and a profit of 500 dollars is the next month, what is my average profit for the two months?

OpenStudy (anonymous):

in other words, what is my average profit per month?

OpenStudy (toxicsugar22):

750

OpenStudy (anonymous):

okay, good. how did you figure that out?

OpenStudy (anonymous):

doing good sugar haha (:

OpenStudy (toxicsugar22):

i added 1000 and 500 and divided by 2 to get the average

OpenStudy (anonymous):

good. so what if I told you that my profit for first month was 1000, second month was 500 and third month was 600. what is my average profit?

OpenStudy (toxicsugar22):

700

OpenStudy (anonymous):

very good. what did you do to get that answer?

OpenStudy (toxicsugar22):

i added all three and divided it by three

OpenStudy (anonymous):

good. so what if I told you my profit for the first month was x and my profit for second month was y, what is my average profit?

OpenStudy (toxicsugar22):

would it be x+y divided by 2

OpenStudy (anonymous):

exactly! very good!

OpenStudy (anonymous):

so your expression is average = (x+y)/2

OpenStudy (anonymous):

if you had 10 numbers, then your average would be (a+b+c+d+e+f+g+h+i+j)/10

OpenStudy (anonymous):

where a,b,c... are the 10 numbers

OpenStudy (toxicsugar22):

ok and can u tell me how do i write the recipocal of the average

OpenStudy (anonymous):

reciprocal of a number x is 1/x. the reciprocal is a number, which when multiplied by x, gives you 1. 1 is called the multiplicative identity. so, when you multiply x times 1/x, you get x/x = 1 similarly, what is the reciprocal of the number (x+y)/2 ?

OpenStudy (toxicsugar22):

ok i got it its 2/(x+y)

OpenStudy (anonymous):

okay very good!

OpenStudy (toxicsugar22):

and one more quesion, for the same numbers x and y, write an expresiion for their average of their recipocals how do i do that

OpenStudy (anonymous):

what is the reciprocal of x? what is the reciprocal of y?

OpenStudy (toxicsugar22):

x/1 and y/1 is that right

OpenStudy (anonymous):

so if you multiply x times x/1 you will get 1? are you sure?

OpenStudy (toxicsugar22):

im srry 1/x and 1/y

OpenStudy (anonymous):

okay good. so now you have two numbers 1/x and 1/y. what is their average?

OpenStudy (toxicsugar22):

2/xy

OpenStudy (anonymous):

no. thats wrong. the average of two numbers a and b is (a+b)/2 so what is the average of two numbers 1/x and 1/y?

OpenStudy (toxicsugar22):

(1/x+1/y)/2

OpenStudy (anonymous):

okay. good. that is your answer. you can simplify the expression further if you want, but that is the answer to your question. 1/x+1/y = (x+y)/xy so (1/x+1/y)/2 = (x+y)/2xy

OpenStudy (toxicsugar22):

ok thank you

OpenStudy (anonymous):

or rather, 2(x+y)/xy

OpenStudy (anonymous):

no, its the first one. sorry.

OpenStudy (toxicsugar22):

ok thanks

OpenStudy (anonymous):

(:

OpenStudy (toxicsugar22):

hey can u help me with another question math mind can u try

OpenStudy (toxicsugar22):

Andy drives 300 miles to Lake Tahoe at 70 miles per hour and returns home at 50 mph. What is his average speed for the round trip. (It is not 6o miles per hour.) Write expressions for the time it takes for each leg of the trip if Andy drives a distance d, at speed r base 1 and returns at speed r base 2

OpenStudy (anonymous):

post it up as a new question, it's better